# Limitations on Safe, Trusted, Artificial General Intelligence

### v1

@@ A.2 Impossibility of solving feasibility when a path exists in the graph
If an AI system $A$ is safe and always halts in some time $T$, then for a fixed constant $c$ and the time limit $T+c$, there is a program, input pair for which there is a short, constant-sized proof that the instance always halts in at most $T+c$ steps, but for a safe AI system $A$ which always halts in time $T$ the probability of $A$ giving the answer ‘always halts in given time limit’ is 0.
@@
If $A$ is safe and always terminates in time $T$, then $\mathtt{self\_Turing\_T\_v2(self\_Turing\_T\_v2)}$ always halts in time at most $T+c$. Moreover, if $A$ is safe then it has 0 probability of giving the answer ‘always halts in given time limit’ on whether $\mathtt{self\_Turing\_T\_v2(self\_Turing\_T\_v2)}$ halts in time at most $T+c$.
@@
Suppose $A$ outputs ‘always halts in given time limit’ on the given input on some randomness. Whenever $A$ outputs ‘always halts in given time limit’, the program enters an infinite loop and never halts. This is a contradiction if $A$ is safe, and hence if $A$ is safe it outputs ‘always halts in given time limit’ with probability 0.
Now, if $A$ does not output ‘always halts in given time limit’ on the input, then the program will enter the else block and immediately halt. Since $A$ runs for at most $T$ steps, the program then halts in time at most $T+c$. Therefore, if $A$ is safe then the program always halts in time at most $T+c$.
A.3 Proof of Theorem 4.2
Theorem 4.2. If the AI system $A$ is well-behaved and calibration-safe for program verification, then there is a program $P$ which provably halts with probability at least $0.99$, but $A$ abstains with probability 1 on the program $P$.
@@
If $A$ is calibration-safe and well-behaved then $\mathtt{G\ddot{o}del\_program\_random}$ halts with probability at least $0.99$, but the calibration-safe AI system $A$ will output ‘don’t know’ with probability 1 on $\mathtt{G\ddot{o}del\_program\_random}$.
@@
    $p\in(0,0.5)$: In this case, with probability at least $0.99$ the best arm identification procedure determines that $\text{arm}_{1}$ is better than $\text{arm}_{2}$. When $\text{arm}_{1}$ is determined to be better than $\text{arm}_{2}$, the program enters the else block in line 9. Therefore, in this case $\mathtt{G\ddot{o}del\_program\_random}$ terminates with probability at least $0.99$. Here too, $A$ is not calibration-safe, since it claims that $\mathtt{G\ddot{o}del\_program\_random}$ terminates with probability $p<0.5$.
In each of these cases, $A$ is not calibration-safe. Therefore, for $A$ to be calibration-safe, we must have $p=0$, and that $A$ outputs ‘don’t know’ with probability 1. If $p=0$, then with probability at least $0.99$ the best arm identification procedure determines that $\text{arm}_{1}$ is better than $\text{arm}_{2}$, and $\mathtt{G\ddot{o}del\_program\_random}$ terminates. Therefore, if $A$ is calibration-safe, then $\mathtt{G\ddot{o}del\_program\_random}$ halts with probability at least $0.99$.

### v2

@@ A.3 Impossibility of solving feasibility when a path exists in the graph
If an AI system $A$ is accurate and always halts in some time $T$, then for a fixed constant $c$ and the time limit $T+c$, there is a program, input pair for which there is a short, constant-sized proof that the instance always halts in at most $T+c$ steps, but for an accurate AI system $A$ which always halts in time $T$ the probability of $A$ giving the answer ‘always halts in given time limit’ is 0.
@@
If $A$ is accurate and always terminates in time $T$, then $\mathtt{self\_Turing\_T\_v2(self\_Turing\_T\_v2)}$ always halts in time at most $T+c$. Moreover, if $A$ is accurate then it has 0 probability of giving the answer ‘always halts in given time limit’ on whether $\mathtt{self\_Turing\_T\_v2(self\_Turing\_T\_v2)}$ halts in time at most $T+c$.
@@
Suppose $A$ outputs ‘always halts in given time limit’ on the given input on some randomness. Whenever $A$ outputs ‘always halts in given time limit’, the program enters an infinite loop and never halts. This is a contradiction if $A$ is accurate, and hence if $A$ is accurate it outputs ‘always halts in given time limit’ with probability 0.
Now, if $A$ does not output ‘always halts in given time limit’ on the input, then the program will enter the else block and immediately halt. Since $A$ runs for at most $T$ steps, the program then halts in time at most $T+c$. Therefore, if $A$ is accurate then the program always halts in time at most $T+c$.
A.4 Proof of Theorem 4.4
See 4.4
Proof.
For any property non-trivial, semantic property $\pi$, let $\texttt{valid\_program\_for\_}\pi$ be some program which has property $\pi$, and $\texttt{invalid\_program\_for\_}\pi$ be some program which does not have property $\pi$. Note that since $\pi$ is a non-trivial property, both these programs exist. Let $A$ be an AI system which takes some program as input, and verifies if the input program has the property $\pi$. Note that $A$ is well-behaved, its execution does not automatically lead to some program having property $\pi$. Now consider Algorithm 6, for any input $I$.
Algorithm 6 Rice_program for property $\pi$
1:procedure $\mathtt{Rice\_program}$(Input II)
[…]
3:   return $\texttt{invalid\_program\_for\_}\pi(I)$
4:  else
5:   return $\texttt{valid\_program\_for\_}\pi(I)$
6:  end if
7:end procedure
We claim that if $A$ is accurate, then it cannot output that Rice_program has property $\pi$. This is true by contradiction, if $A$ outputs that Rice_program has property $\pi$, then Rice_program calls $\texttt{invalid\_program\_for\_}\pi$, and hence does not have property $\pi$. Note that if $A$ does not output that Rice_program has property $\pi$, then Rice_program always calls $\texttt{valid\_program\_for\_}\pi$, and hence Rice_program has property $\pi$.
Therefore if $A$ is accurate, then Rice_program has property $\pi$, but $A$ cannot output that Rice_program has property $\pi$. Hence an accurate and trusted system for verifying property $\pi$ cannot be a human-level reasoning system for the task.
A.5 Proof of Theorem 5.2
See 5.2
@@
If $A$ is calibrated and well-behaved then $\mathtt{G\ddot{o}del\_program\_random}$ halts with probability at least $0.99$, but the calibrated AI system $A$ will output ‘don’t know’ with probability 1 on $\mathtt{G\ddot{o}del\_program\_random}$.
@@
    $p\in(0,0.5)$: In this case, with probability at least $0.99$ the best arm identification procedure determines that $\text{arm}_{1}$ is better than $\text{arm}_{2}$. When $\text{arm}_{1}$ is determined to be better than $\text{arm}_{2}$, the program enters the else block in line 9. Therefore, in this case $\mathtt{G\ddot{o}del\_program\_random}$ terminates with probability at least $0.99$. Here too, $A$ is not calibrated, since it claims that $\mathtt{G\ddot{o}del\_program\_random}$ terminates with probability $p<0.5$.
In each of these cases, $A$ is not calibrated. Therefore, for $A$ to be calibrated, we must have $p=0$, and that $A$ outputs ‘don’t know’ with probability 1. If $p=0$, then with probability at least $0.99$ the best arm identification procedure determines that $\text{arm}_{1}$ is better than $\text{arm}_{2}$, and $\mathtt{G\ddot{o}del\_program\_random}$ terminates. Therefore, if $A$ is calibrated, then $\mathtt{G\ddot{o}del\_program\_random}$ halts with probability at least $0.99$.